### v2

@@ 4.3. RQ3: What phases of the development process are supported by the DSLs?
Only 2 DSLs express (business) process models with a visual representation of security considerations. While Rodríguez et al. (Rodríguez et al., 2007) extend the BPMN, the Security Policy Model (Wolter et al., 2008; Menzel et al., 2009; Wolter et al., 2009) defines its own process model notation.
@@
With 15 DSLs (13 % of all DSLs), many DSLs are used to create design models that express a secure design of the planned system. For example, as shown in Fig. 8, UMLsec is used to plan secure deployments of software systems.
@@
Corresponding to the SDLC phase of coding, where most security DSLs are used, with 32 DSLs (27 %), code is the second-largest artifact category. DSLs are used to create custom code artifacts, such as STATL (see LABEL:lst:statl) or Cryptol (see LABEL:lst:cryptol), as well as code artifacts in common general purpose languages, such as TASTYL which is based on python. We recall that only 17 % of the DSLs facilitate security with a scope limited to software, for example, by addressing side-channel security or memory-safety. Therefore, many DSLs with code artifacts feature defense mechanisms that consider the interaction of several system components, e.g., distributed systems.
@@
Hardware description languages that express security considerations of hardware, typically in the form of logic circuits that can be generated and come with some security guarantees, are provided by 5 DSLs (4 %).
@@
Attacks and, in the case of some DSLs, appropriate defenses can be described using 11 of the DSLs (9 %). Most of these DSLs use a graph-based representation such as attack trees (Kordy et al., 2010) or probabilistic attack-defense graphs (Katsikeas et al., 2019; Johnson et al., 2018; Widel et al., 2023).
@@
With 43 DSLs (36 %), policy files are the largest artifact category we encountered. Since policies are a fundamental aspect of security and a component of various defense mechanisms, especially access control systems, which were the most common category of security aspects among all security DSLs (see Section 4.2), this is an expected observation. Within this category, 29 DSLs are used to specify plain policies, while 13 DSLs (11 %) are used to create concrete access control models that go beyond simple policies. In addition, PPLv2 (Cerbo et al., 2015) specifies detailed models for access and usage control, which even goes beyond what can be expressed in the policy files or access control models considered above.
@@
Some form of intrusion detection rules can be defined using 3 of the DSLs (3 %). AIDD (Souissi, 2015) and IDS-DL (Amine et al., 2019) facilitate the specification of intrusion detection rules, which could be seen as special policies or attack and defense descriptions. Haka (Arkoon Network Security and ParisTech, 2014) allows specifying communication protocols in combination with filtering rules to identify and address malware network activities.
@@
Lastly, FAL (Zawoad et al., 2013) is used to specify configuration files of a secure logger.
Integration with other artifacts
@@
First, 4 security DSLs (3 %) enrich requirements expressed in different artifacts with security-related aspects such as relevant threats for concrete requirements.
@@
Similarly, 20 DSLs (17 %) enrich design models with relevant security objectives, threats, and planned security features. For example, in Figure 8, the UMLsec security DSL is used to enrich a UML deployment diagram with information on the criticality of communicated information and requirements on the security of used communication paths.
@@
While 13 DSLs (11 %) are integrated with policies, thereof, only 9 DSLs express themselves policies or access/usage control models. Usually, these are integrated or compatible with a policy standard, such as XACML (6 DSLs).
@@
Overall, 10 DSLs (9 %) are integrated with runtime components. Mostly, these DSLs are integrated with components of the system at runtime (8 DSLs), usually meaning that the instance of the DSL is executed at runtime in combination with this runtime component, e.g., compiled functional source code. In addition, 2 DSLs are integrated with intrusion detection systems and build the basis for monitoring a system.
@@
Only STREAMS (Kwon et al., 2017) integrates with existing hardware description artifacts by providing a hardware module for use in Verilog. All other DSLs that target hardware description specify these in their own languages, but do not integrate with existing specifications.
@@
Kind of integration
Less than half of the DSLs examined (51 DSLs) are relevant to multiple phases of the SDLC as part of the development of secure systems. We identified five variants in which security DSLs can be integrated into development. It is possible that a single DSL can be integrated into the SDLC using multiple of these variants.
@@
Most of the security DSLs (23 %) serve as languages for planning artifacts. The use of DSLs enables the documentation of security requirements, planned countermeasures, and other security aspects that must be considered or realized in subsequent phases of the SDLC. These serve as requirements for subsequent phases, with the manual usage of artifacts created using the corresponding DSLs representing one of their primary purposes.
@@
Many artifacts created using security DSLs (13 %) are reused in later phases of the SDLC by other security engineering techniques, such as access control policies derived at design time that serve as input to an access control engine at runtime.
@@
Finally, the Access Control Framework (Pavlich-Mariscal et al., 2010) also allows tracing security requirements across different artifacts created in the SDLC.
@@
The vast majority of the security DSLs (58 %) has been designed with a single artifact in mind and is used in isolation.
Observation 3. Security DSLs are used in all phases of the software development life cycle (SDLC), but 58 % of the DSLs do not feature any means of integration into the SDLC. There is a significant focus on design-time and code-level DSLs. Particularly, deployment and testing are underrepresented. The most common artifacts of security DSLs are security policies, code, or design models. While security DSLs sometimes are used in multiple SDLC phases, this is primarily the case for security DSLs targeting the requirements engineering or design phase.
4.4. RQ4: What types of DSLs exist, and how are they used?
For embedding security aspects into development processes as discussed above, DSLs can be specified following various language design principles. To get a better understanding of the technical perspective on security DSLs, first, we look at how the security DSLs are defined and what are their characteristics. To this end, we discuss whether the DSLs are an extension of an existing language or are newly designed languages and how instances are represented. Thereafter, we discuss how developers work with the DSLs and how they are supported doing so.
Language design
When examining the artifacts created by the security DSLs, we observed that some of them are defined entirely by syntax, while others are enriched versions of other languages, such as extended UML models. In general, language engineering typically distinguishes between two types of language design, both of which we observed in the security DSLs we studied.
@@ 4.4. RQ4: What types of DSLs exist, and how are they used?
Unlike external DSLs, internal DSLs, which make up the remaining 34 % of DSLs analyzed, enrich an existing language and use its language infrastructure. Among the investigated internal DSLs, the most common host language is the Unified Modeling Language (UML), which is extended by 40 % of all internal security DSLs via so-called UML profiles. In total, 50 % of internal security DSLs extend modeling languages, such as SysML, BPMN, and AADL. The second-largest group of internal DSLs extend programming languages (35 %). Thereby, Java and Haskel are the most extended languages (4 DSLs each), followed by ASML, F*, F#, Javascript, Promela, and Scala with one DSL each. Finally, three DSLs each enrich policy languages (IF, WS-Policy, and XCAML) and requirements languages (GRL and two times SI*). For example, Asnar et al. (Asnar et al., 2011) extend the SI* goal modeling language (Massacci et al., 2010) with threat modeling capabilities.
Relation to existing DSLs
The DSLs analyzed are usually new languages, which in the case of internal DSLs may be related to an underlying modeling or programming language. However, 18 % of them are based on an existing DSL and extend it with additional aspects. Extended security DSLs are mainly DSLs for specifying access control policies (31 %), e.g., WS-CoL and WS-Policy4MASC extend the WS-Policy DSL with additional policy concepts. Secure code DSLs are also being extended, e.g., Fabric and $J_{E}$ both extend Jif with additional concepts. Furthermore, with 35 %, the percentage of threat modeling DSLs which extend another security specific language is remarkably high.
Instance representation
The second characteristic of a language is how instances are represented and made accessible to users. This can be entirely textual, visual, or a combination of the two.
@@
The majority of DSLs (69 %) are textual languages. The DSLs define a textual syntax for the language, allowing users to specify instances in any text editor. DSLs such as Jif extend the language syntax of Java with additional elements and can be used in a similar way to standard Java source code. STARTL, Paragon, Cryptol, and many others define their own programming language-like syntax. Another style of textual syntax is used by WS-Policy, the DSLs that extend it, and many other policy languages. These often use an XML syntax. Almost all DSLs in the information flow control and cryptography categories are textual languages.
@@
STS and UMLsec are examples for a visual DSLs. Figure 6 provides a partial example of a social diagram in STS. The visual notation of STS is oriented on SI* goal models, while UMLsec extends the UML. In the back, both models are serialized as XML files, but these are not intended for manual use. Overall, 24 % of the DSLs we examined work on a visual representation.
@@
Interestingly, 64 % of problem-oriented DSLs are visual, and 11 % are textual and visual.
Editing support
To allow users to work efficiently and effectively with DSLs it is important to provide them with editing support. Editing support ranges from simple batch syntax checking, such as usually provided by compilers, over its embedding into IDEs for live feedback on the syntax, and syntax highlighting to semantic feedback on the expressed aspects or suggestions for auto-completion. Among the DSLs, we identified four degrees of editing support (cf. Fig. 9).
[…]
Figure 9. Statistics on editing support and instance representation and language design (from left to right).
@@
Most DSLs come with no editing support at all (64 %). Users have to write instances with any editor and are not supported in doing so. Mistakes only get recognizable, in the case of errors during compilation, interpretation, or other forms of DSL usage.
@@
The highest degree of editing support comprises instance-specific checks of the expressed semantics and sophisticated editing features such as auto-completion and is provided by 13 % of the DSLs. This degree of editing support is mainly implemented via language-engineering frameworks such as xtext that automatically generate editing support features or UML-based languages that are enriched with semantic checks embedded via OCL constraints.
@@
4.5. RQ5: What are the semantics of the DSLs?
@@
Figure 10. Semantic usage of the DSLs (one DSL can be used in multiple ways).
@@
The majority of DSLs (52 DSLs, or 43 % of all DSLs) provide support to developers by analyzing the software system as represented by an instance of the DSL for security issues, such as for identifying insecure design decisions such as the lack of encrypted communication of sensitive data that should be addressed by software engineers.
@@
There are 24 DSLs (20 %) that are intended to be the source of an automated transformation into another instance format, i.e. another DSL. Typically, these DSLs allow for a simpler specification of the security aspects expressed than the usually more generic target format and domain-specific knowledge that is hidden by the DSL is added through the transformation. In most cases, access control policies derived at design time are transformed into other formats that can be executed at runtime, e.g., by an access control engine.
@@
The compilation of an DSL instance into an executable format in machine code is supported by 20 DSLs (17 %). These DSLs assemble executable programs after compilation, e.g. Jif is an extension of the Java programming language with information flow constraints that compiles to Java bytecode can be executed on the JVM.
@@
Lastly, 33 DSLs (28 %) are executed by an interpreter. Unlike compiled DSLs, no executable machine code is produced, but the DSL is executed as it is by the interpreter. DSLs used in this way mainly cover security constraint specifications such as access control policies that are enforced by a runtime engine.
Automation
Security DSLs are commonly provided with tooling that automates certain aspects of their semantic usage. Among the DSLs that come with such tooling, the average number of automated semantic usages is 1.2, which is much lower than the intended average of 1.5 usages.
@@
While the semantic usage of capturing knowledge for the use by humans is the only semantic usage of security DSLs that cannot be automated and only 15 DSLs have solely this purpose, 33 DSLs (28 %) do not come with any automation. Often, some automation is outlined in future work but has never been realized.
@@
The majority of security DSLs (36 %) come with an automation of their proposed analyses, e.g., for identifying security issues. While 52 DSLs propose analysis as semantic usage, with 43 DSLs most of them actually automate this usage (83 %). Analyzes are mainly hard-coded but are also frequently implemented using query or constraint languages such as the OCL.
@@
All security DSLs that propose a transformation, with the exception of four (83 % of the DSLs that propose transformations), automate this transformation. In total, 20 DSLs (17 %) automate transformations. Transformations are mostly implemented in general purpose languages and model transformation languages build the exception.
@@
All 9 DSLs that propose code generation actually implement tooling for generating source code from the DSL instances.
@@
Only 9 of the 20 of the DSLs that propose to compile the instances into an executable format actually implement such a compiler (60 %).
@@
Similar to compilers also for only 21 of the DSLs (64 %) for which an interpretation at runtime is proposed such an interpreter is actually implemented.
When it comes to the actual analysis of security DSLs for detecting vulnerabilities or integrating them with other SDLC phases through the transformation into another format, 83 % of the DSLs come with such tooling. However, putting the DSLs actually into action remains mainly on a theoretical basis and is often not implemented (only for 60 % of compilation and 64 % of interpretation). One exception is code generation for usage by developers as part of their implementations which is implemented for all 9 DSLs for which such a code generation has been proposed.
@@
4.6. RQ6: How effective are the DSLs?
In our last research question, we focus on the effectiveness of the security DSLs. First we look at how the effectiveness of the security DSLs is assessed. Here we observed both formal proofs of concrete security guarantees and empirical evaluations. While 3 security DSLs do not come with any of the two, one of them is an open standard and widely applied in practice. Therefore, we also investigated whether the security DSLs have been applied in practice, e.g., as part of a case study.
Guarantees and formal proofs
Overall, 40 security DSLs (33 %) state some form of security guarantee. However, there is a strong correlation between the availability of an attacker model and a security guarantee. We observed that if an attacker model is available, 79 % of these DSLs have a security guarantee and explicitly state the security objective. On the other hand, DSLs with a focus on Access Control do not have an attacker model (81 %) and do not provide a security guarantee (93 %) in most cases. If a security guarantee is provided by a security DSL, it should be proven that this guarantee holds under any conditions. Therefore, we investigated whether security DSLs are accompanied by formal proofs.
Not applicable.:
The majority of the security DSLs do not come with a formal proof since these are not applicable to them, e.g., due to their informal nature. Only 33 % of the DSLs provide a security guarantee, for example, many languages in the field of information flow control assert a form of non-interference, a concept introduced by Goguen and Meseguer (Goguen and Meseguer, 1982), which implies that the observable output of a system must be independent of its sensitive input. For the other 67 %, formal proofs are not applicable due to the lack of a security guarantee to provide a proof for.
Formal proof.:
Out of the set of DSLs with security guarantee, 73 % provide a formal proof that assures the correctness of the claimed guarantee. These are mostly from domains with strong formal foundations, e.g., cryptography or information flow control.
No proof.:
Unfortunately, 12 DSLs remain that contain security aspects that should be accompanied by a formal proof but for which no formal proof is provided.
Empirical evaluation
To get a better understanding how the security DSLs are evaluated, we first assessed using which empirical method the DSLs as well as supporting tooling are evaluated.

### v3

@@ 4.3.2. Artifacts Specified Using the DSLs
Two DSLs express (business) process models with a visual representation of security considerations. While Rodríguez et al. (257) extend the BPMN, the Security Policy Model (256; 259; 293) defines its own process model notation.
@@
With 15 DSLs (13 %), many DSLs are used to create design models of the planned system. Fig. 8 shows how UMLsec is used to plan secure deployments of software systems.
@@
Corresponding to the SDLC phase of coding, where most DSLs are used, with 32 DSLs (27 %), code is the second-largest artifact category. DSLs are used to create custom code artifacts, such as STATL (see LABEL:lst:statl), as well as code artifacts in common general purpose languages, such as TASTYL which is based on python. We recall that only 17 % of the DSLs facilitate security with a scope limited to software, for example, by addressing side-channel security or memory-safety. Therefore, many DSLs with code artifacts feature defense mechanisms that consider the interaction of several system components, e.g., distributed systems.
@@
Five DSLs (4 %) are hardware description languages that express security considerations of hardware, typically in the form of logic circuits that can be generated and come with some security guarantees.
@@
Eleven DSLs (9 %) are used to describe attacks and some of them also appropriate defenses. Most of these DSLs use a graph-based representation such as attack trees (340) or probabilistic attack-defense graphs (17; 22; 189).
@@
With (43 DSLs, 36 %), policy files are the largest artifact category we encountered. Since policies are a fundamental aspect of security and a component of various defense mechanisms, especially access control systems, which were the most common category of security aspects among all DSLs (see Sec. 4.2), this is an expected observation. Within this category, 29 DSLs are used to specify plain policies, while 13 DSLs (11 %) are used to create concrete access control models that go beyond simple policies. One DSL (PPLv2(118)) specifies detailed models for access and usage control, which even goes beyond what can be expressed in the policy files or access control models considered above.
@@
Three DSLs (3 %) define some form of intrusion detection rules. AIDD (69) and IDS-DL(28) facilitate the specification of intrusion detection rules, which could be seen as special policies or attack and defense descriptions. Haka(266) allows specifying communication protocols in combination with filtering rules to identify and address malware network activities.
@@
Lastly, FAL(222) is used to specify configuration files of a secure logger.
Following the low coverage of the test and deployment phases, the artifacts expressed using DSLs do not cover test cases or concrete deployment artifacts. Security configurations are the closest related artifacts to deployment, but only one DSL considers them. Test-specific artifacts that are explicitly expressed using a DSL are entirely missing.
4.3.3. Integration With Other Artifacts
@@
Four DSLs (3 %) enrich requirements expressed in different artifacts with security-related aspects such as relevant threats for concrete requirements.
@@
Similarly, 20 DSLs (17 %) enrich design models with security objectives, threats, and planned features. For example, in Figure 8, UMLsec is used to enrich a UML deployment diagram with information on data sensibility and the security requirements of its communication.
@@
Thirteen DSLs (11 %) are integrated with policies, thereof, nine DSLs express themselves policies or access/usage control models. Usually, these are integrated or compatible with a policy standard, such as XACML (6 DSLs).
@@
Nine percent of the DSLs are integrated with runtime components. Eight DSLs are integrated with components of the system at runtime, usually meaning that the instance of the DSL is executed at runtime in combination with this runtime component, e.g., compiled functional source code. Two DSLs are integrated with intrusion detection systems.
@@
Only STREAMS(43) integrates with existing hardware description artifacts by providing a hardware module for use in Verilog. All other DSLs that target hardware description specify these in their own languages, but do not integrate with existing specifications.
@@
4.3.4. Kind of Integration
Fewer than half of the DSLs (51 DSLs) are relevant to multiple development phases. We identified five variants in which DSLs can be integrated into the development. A single DSL can be integrated using multiple of these variants.
@@ 4.3.4. Kind of Integration
Most of the DSLs (23 %) serve as languages for planning artifacts. The use of DSLs enables the documentation of security requirements, planned countermeasures, and other security aspects that must be considered or realized in subsequent phases of the SDLC. These serve as requirements for subsequent phases, with the manual usage of artifacts created using the corresponding DSLs representing one of their primary purposes.
@@
Many artifacts created using DSLs (13 %) are reused in later phases of the SDLC by other security engineering techniques, such as access control policies derived at design time that serve as input to an access control engine at runtime.
@@
Finally, the Access Control Framework (269) also allows tracing security requirements across different artifacts created in the SDLC.
@@
Most DSLs (58 %) have been designed with a single artifact and are used in isolation.
Observation 3. Security DSLs are used in all phases of the SDLC, but 58 % of them do not feature any means of integration into the SDLC. There is a significant focus on design-time and code-level DSLs, while deployment and testing are underrepresented. The most common artifacts are security policies, code, or design models. Primarily DSLs targeting the requirements engineering or design phase are used in multiple SDLC phases.
4.4. RQ4: What Types of Security DSLs Exist, and How are They Used?
For embedding security aspects into development processes as discussed above, DSLs can be specified following various language design principles. To get a better understanding of the technical perspective on security DSLs, first, we look at how the DSLs are defined and what are their characteristics. To this end, we discuss whether the DSLs are an extension of an existing language or are newly designed languages and how instances are represented. Thereafter, we discuss how developers work with the DSLs and how they are supported.
4.4.1. Language Design
When examining the artifacts created using the DSLs, we observed that some are defined entirely by the DSL’s syntax, while others are enriched versions of other languages.
@@ 4.4.1. Language Design
Internal DSLs (34 %) enrich an existing language and use its language infrastructure. The most common host language we observed is the Unified Modeling Language (UML), which is extended by 40 % of all internal DSLs via so-called UML profiles. In total, 50 % of internal DSLs extend modeling languages, such as SysML, BPMN, and AADL. The second-largest group of internal DSLs extend programming languages (35 %). Thereby, Java and Haskel are the most extended languages (4 DSLs each). Finally, three DSLs each enrich policy languages (IF, WS-Policy, and XCAML) and requirements languages (GRL and two times SI*). For example, Asnar et al. (233) extend the SI* goal modeling language (Massacci et al., 2010) with threat modeling capabilities.
4.4.2. Relation to Existing DSLs
The DSLs analyzed are usually new languages, which in the case of internal DSLs may be related to an underlying modeling or programming language. However, 18 % of them are based on an existing DSL and extend it with additional aspects. Extended DSLs are mainly DSLs for specifying access control policies (31 %), e.g., WS-CoL and WS-Policy4MASC extend the WS-Policy DSL with additional policy concepts. Secure code DSLs are also being extended, e.g., Fabric and $J_{E}$ both extend Jif with additional concepts. In addition, the percentage of threat modeling DSLs that extend another security-specific language is remarkably high at 35 %.
4.4.3. Instance Representation
The second characteristic of a language is how instances are represented and made accessible to users. This can be entirely textual, visual, or a combination.
@@ 4.4.3. Instance Representation
The majority of DSLs (69 %) are textual languages. The DSLs define a textual syntax for the language, allowing users to specify instances in any text editor. DSLs such as Jif extend the language syntax of Java with additional elements and can be used in a similar way to standard Java source code. STARTL, Paragon, Cryptol, and many others define their own programming language-like syntax. Policy languages often use an XML syntax. Almost all DSLs in the information flow control and cryptography categories are textual languages.
@@
STS and UMLsec are examples for a visual DSLs. Figure 6 provides a partial example of a social diagram in STS. The visual notation of STS is oriented on SI* goal models, while UMLsec extends the UML. In the back, both models are serialized as XML files, but these are not intended for manual use. Overall, 24 % of the DSLs we examined work on a visual representation. Interestingly, 64 % of problem-oriented DSLs are visual, and 11 % are textual and visual.
@@
4.4.4. Editing Support
To allow users to work efficiently and effectively with DSLs it is important to provide them with editing support (Barros et al., 2022). Editing support ranges from simple batch syntax checking, such as usually provided by compilers, over its embedding into IDEs for live feedback on the syntax, and syntax highlighting to semantic feedback on the expressed aspects or suggestions for auto-completion. Among the DSLs, we identified four degrees of editing support (cf. Fig. 9).
None64%Basic16%Extended8%Sophisticated13%Textual69%Visual24%Text. & Vis.7%External66%Internal34%
Figure 9. Statistics on editing support and instance representation and language design (from left to right)
@@ 4.4.4. Editing Support
Most DSLs come with no editing support at all (64 %). Users have to write instances with any editor and are not supported in doing so.
@@
The highest degree of editing support comprises instance-specific checks of the expressed semantics and sophisticated features such as auto-completion and is provided by 13 % of the DSLs. This degree is mainly implemented via language-engineering frameworks, such as Xtext (Bettini, 2016) that automatically generate editing support features (Barros et al., 2022) or UML-based languages that are enriched with semantic checks embedded via OCL constraints.
@@
4.5. RQ5: What are the Semantics of the Security DSLs?
@@ 4.5.1. Usage Purpose
055101015152020252530303535404045455050KnowledgeAnalyzeTransformCodegenCompileInterpretNumber of DSLs
Figure 10. Semantic usage of the DSLs (one DSL can be used in multiple ways)
@@
52 DSLs (43 %) provide support to developers by analyzing the system as represented by an instance for security issues, such as identifying insecure design decisions.
@@
There are 24 DSLs (20 %) that are intended to be the source of an automated transformation into another instance format, i.e., another DSL. Typically, these DSLs allow for a simpler specification of the security aspects expressed than the usually more generic target format and domain-specific knowledge that is hidden by the DSL is added through the transformation. In most cases, access control policies derived at design time are transformed into other formats that can be executed at runtime, e.g., by an access control engine.
@@
20 DSLs (17 %) compile an instance into an executable format.
@@
33 DSLs (28 %) are executed by an interpreter. DSLs used in this way mainly cover security constraints such as access control policies that are enforced by a runtime engine.
4.5.2. Automation
DSLs are commonly provided with tooling that automates certain aspects of their semantic usage. Among the DSLs that come with such tooling, the average number of automated semantic usages is 1.2, which is much lower than the intended average of 1.5 usages.
@@ 4.5.2. Automation
While capturing knowledge for the use by humans is the only semantic usage purpose of DSLs that cannot be automated and only 15 DSLs have solely this purpose, 33 DSLs (28 %) do not come with any automation. Tough, some automation is often outlined as future work.
@@
The majority of DSLs (36 %) come with an automation of their proposed analyses, e.g., for identifying security issues. While 52 DSLs propose analysis as semantic usage, with 43 DSLs most of them actually automate this usage (83 %). Analyzes are mainly hard-coded, but are also frequently implemented using query or constraint languages such as the OCL. While automated analyses mostly do not provide guaranteed security aspects, they help in identifying vulnerabilities and increase the trust in the artifacts expressed using the DSL.
@@
All DSLs that propose a transformation, with the exception of four (83 % of the DSLs that propose transformations), automate this transformation. In total, 20 DSLs (17 %) automate transformations. Transformations are mostly implemented in general purpose languages and model transformation languages build the exception.
@@
All nine DSLs that propose code generation actually implement tooling for generating source code from the DSL instances. This way, vulnerabilities can be avoided and the security of the system can be increased.
@@
Only 9 of the 20 DSLs that propose to compilation actually implement a compiler (60 %).
@@
Similarly, only 64 % of the 20 DSLs that propose interpretation also realize it.
83 % of the DSLs come with tooling for automated vulnerability detection or for integration with other SDLC phases through transformation. However, putting the DSLs actually into action remains mainly theoretical and is often not implemented. One exception is code generation for usage by developers as part of their implementations which is implemented for all 9 DSLs for which such a code generation has been proposed.
@@
4.6. How are the Security DSLs Evaluated?
Our final research question focuses on the evaluation of the DSLs, in particular the evaluation methodology and objective. Three DSLs do not come with any evaluation, but one of them is an open standard and widely used in practice. Therefore, we also investigated whether the DSLs have been applied in practice, e.g., as part of a case study.
4.6.1. Empirical Methods
Empirical evaluations assess the claims associated with the DSLs by gathering and analyzing observable evidence or data. Claims can relate to the security goal (e.g., effectiveness to prevent security issues) or quality aspects (e.g., usability). To obtain an overview, we first determined using which empirical method the DSLs and their supporting tooling are evaluated.